OpenStudy (anonymous):

Solve for x: -2(x + 3) = -2(x + 1) - 4.

OpenStudy (anonymous):

-2x - 6 = -2x - 2 - 4 -2x - 6 = -2x -6 if you move over the -2x to the left side of the equality sign you'll get 0x = 0 So no solution

OpenStudy (cggurumanjunath):



OpenStudy (anonymous):

thanks, thats what i got. just making sure

OpenStudy (anonymous):

Yep :)

OpenStudy (anonymous):

Couple more ? Solve for x: 3(x + 1) = -2(x - 1) + 6 Solve for x: 3x - 5 = 2x + 6. Solve for x: the quantity of x plus 16 all over 3 = 3x @jadeybells ?;)

OpenStudy (anonymous):

1) 3x + 3 = -2x + 2 + 6 3x + 3 = -2x + 8 +2x +2x 5x + 3 = 8 -3 -3 5x = 5 x = 1

OpenStudy (anonymous):

3x - 5 = 2x + 6. -2x -2x x - 5 = 6 +5 +5 x = 11

OpenStudy (anonymous):

(x+16)/3 = 3x x+16 = 9x 9x - x = 16 8x = 16 x = 16/8 x = 2

OpenStudy (anonymous):

Thanks (; , i greatly appreciate it ^_^

OpenStudy (anonymous):

Yuppers! I enjoy helping with math

OpenStudy (anonymous):

Youre pretty good at it .

OpenStudy (anonymous):

Average, to be honest :/ Lol

OpenStudy (anonymous):

haha, nahh youre well above it , Facebook ? you know, for anymore questions ?;)

OpenStudy (anonymous):

Lol I have a Google, I don't use my facebook anymore.

OpenStudy (anonymous):

Ah, i don't have one of those. /:

OpenStudy (anonymous):

Make one, it's pretty simple

OpenStudy (anonymous):

jk, don't actually make one /)o^

OpenStudy (anonymous):

https://plus.google.com/u/0/101315281650224147628/posts too late lol

OpenStudy (anonymous):

Just added ya :)

OpenStudy (anonymous):

okk

OpenStudy (anonymous):

:)